OpenStudy (anonymous):

A bag contains colored marbles as shown in the table below. Two marbles are drawn at random from the bag. Find each probability. Color Number red 5 green 3 blue 2

OpenStudy (anonymous):

13. P(red, then green) if replacement occurs 14. P(red, then green) if no replacement occurs 15. P(2 red) if no replacement occurs 16. P(2 red) if replacement occurs

OpenStudy (anonymous):

do with and without replacement,ty

OpenStudy (anonymous):

13. P(red, then green) if replacement occurs =5/10*3/10=15/100=3/20 14. P(red, then green) if no replacement occurs 5/10*3/9=15/90=1/6 15. P(2 red) if no replacement occurs =5/10 *4/9=20/90=2/9 16. P(2 red) if replacement occurs =5/10*5/10=25/100=1/4